OpenStudy (geometry_hater):

Someone please teach me how to do this 6 to the 2x power = 21

OpenStudy (anonymous):

using logarithms to find x because there is a property of logarithms that says that log(2^x) = xlog(2), so that is how we get x out of the exponent, but we have to take the log of both sides log(6^2x) = log(21) 2xlog(6) = log(21) divide by log 6 to get 2x by itself 2x = log(21)/log(6) then divide that by two to get x

OpenStudy (geometry_hater):

uhh ok

OpenStudy (anonymous):

? Did what i say help at all?

OpenStudy (geometry_hater):

yeah i have to use some change of base formula to get the answer or something

OpenStudy (anonymous):

just use logs, its pretty quick to do